# Large-Scale Multi-dimensional Knowledge Profiling of Scientific Literature

### v1

@@ 6 Conclusion
  • Sec. A. $ResearchDB$ Source and Data Schema.
@@
Appendix A $ResearchDB$ Source and Data Schema
@@
Database Source and Distribution. The $ResearchDB$ comprises over 100,000 papers from 22 top-tier conferences and journals published between 2020 and 2025. This broad selection ensures comprehensive coverage across key AI and Computer Science sub-domains, as shown in Tab. A1 and Tab. A3.
@@
Schema Documentation. The ResearchDB schema captures five main dimensions (Summary, Technical, Analysis, System, Post-Processed) extracted from full paper texts using LLM-assisted parsing. Tab. A2 provides the detailed documentation for the key fields in our structured data records.
Table A2: Revised documentation for the key structural fields in the ResearchDB schema, reflecting the specific output of the LLM-assisted parsing pipeline.
Field Name Data Type Source Purpose/Description
conference/year String/Int Metadata Publication venue and year.
Table A3: Detailed annual paper distribution for the 22 major venues in ResearchDBResearchDB from 2020 to 2025.
Venue Group 2020 2021 2022 2023 2024 2025 Total
Computer Vision (CVPR, ICCV, ECCV) 1358 3272 3715 4509 4868 2871 20593
Systems/Security (MLSYS, OSDI, NDSS, etc.) 216 374 468 467 479 36 2040
Overall Total (22 Venues) 7580 15324 16951 20944 25538 18145 104482
@@
LLM Knowledge Parsing Prompts. We utilized the Deepseek-R1-32B model for semantic parsing of full paper texts to populate the $ResearchDB$ schema (as documented in Appendix A). The parsing fidelity hinges on a carefully designed, JSON constrained prompt template. Below is the primary prompt template used for extracting core technical information (e.g., methods, architecture, training_setup).
@@
13. **Gpu_info**:Obtain the gpu resources used in the article, <total_gpu>*<GPU_MODEL>*<training_time>.
@@
ResearchDB Topic Clustering Configuration. The multidimensional topic structure of $ResearchDB$ was generated through a robust embedding-based pipeline. We first transformed the paper abstracts and key textual segments into semantic vectors using the all-MiniLM-L6-v2 model. These high dimensional embeddings were then processed by UMAP for noise reduction and projection into a lower dimensional space, specifically configured with n_components=40n\_components=40, n_neighbors=120n\_neighbors=120, and min_dist=0.08min\_dist=0.08 (with a fixed random_state=42random\_state=42 for reproducibility). UMAP utilized the cosine metric, which aligns with the semantic nature of the embeddings. Clustering was performed using HDBSCAN, which is optimized for identifying clusters of varying densities; the algorithm was set with a significantly smaller min_cluster_size=50min\_cluster\_size=50 and min_samples=1min\_samples=1, employing the eom (Excess of Mass) cluster selection method. Topic representation was established using C-TF-IDF, configured to filter vocabulary using $\text{MAX\_DF\_PERCENT}=0.9$ and a $\text{MIN\_DF\_ABS\_FLOOR}=50$, utilizing an n-gram range of (1, 2). The full hierarchical structure, derived from the HDBSCAN tree and refined by an external LLM (ChatGPT-5) for descriptive naming, resulted in the fine-grained multi-level taxonomy of over 300 topics presented in Appendix C.
Table A4: Sample list of fine-grained topics extracted from $ResearchDB$, illustrating the specificity and coverage of the taxonomy.
@@
Core Machine Learning 042 Causal Inference for Robustness and Generalization do-calculus, invariant prediction, domain generalization
@@
Hierarchical Topic Tree Structure. A key advantage of the HDBSCAN algorithm is its inherent ability to produce a Hierarchical Structure. We leverage this structure as an organizational framework, providing a non-strict view of clustering relationships above the 324 fine-grained topics. This tree structure is the basis for our hierarchical retrieval system, allowing users to navigate from macro-level concepts (near the root) down to specific fine-grained topics (at the leaves), as shown in Fig. A1.
@@
Topic Naming and Examples. The naming of each fine-grained topic underwent strict quality control: C-TF-IDF was used to extract representative keywords, which were then analyzed and refined by ChatGPT-5 based on paper abstracts within the cluster. This process generated highly semantically clear Topic Names (e.g., 140_Approximation and Fair Clustering Algorithms). Every paper in $ResearchDB$ is uniquely tagged with such a topic name. The table below presents a sample of the most representative fine-grained topics. We have selected topics with ids ranging from 1 to 20 and listed them as Tab. A5
Table A5: Twenty Sample Fine-Grained Topics from $ResearchDB$, Illustrating the Taxonomy’s Scope and Detail.
@@
Evaluation Results. Overall validation demonstrates that the $ResearchDB$ retrieval system, driven by multidimensional knowledge profiling, significantly outperforms traditional keyword or embedding-only retrieval baselines on complex, multi-faceted queries. To illustrate this advantage, we conducted a qualitative assessment contrasting the $ResearchDB$ output with a ChatGPT-5 baseline on a complex query: “How to achieve real-time, high-fidelity speech-to-gesture generation using decoupled diffusion models.” The $ResearchDB$-supported answer (akin to an Academic Review) showed clear superiority in Fig. A2:

### v2

@@ Conclusion
  • Sec. A. $ResearchDB$ Source .
@@
Appendix A $ResearchDB$ Source
@@
Database Source and Distribution. The $ResearchDB$ comprises over 100,000 papers from 22 top-tier conferences and journals published between 2020 and 2025. This broad selection ensures comprehensive coverage across key AI and Computer Science sub-domains, as shown in Tab.˜A1.
@@
LLM Knowledge Parsing Prompts. We utilized the Deepseek-R1-32B model for semantic parsing of full paper texts to populate the $ResearchDB$ schema (as documented in Appendix˜A). The parsing fidelity hinges on a carefully designed, JSON constrained prompt template. Below is the primary prompt template used for extracting core technical information (e.g., methods, architecture, training_setup).
@@
13. **Gpu_info**:Obtain the gpu resources used in the article, <total\_gpu>*<GPU_MODEL>*<training_time>.
@@
ResearchDB Topic Clustering Configuration. The multidimensional topic structure of $ResearchDB$ was generated through a robust embedding-based pipeline. We first transformed the paper abstracts and key textual segments into semantic vectors using the all-MiniLM-L6-v2 model. These high dimensional embeddings were then processed by UMAP for noise reduction and projection into a lower dimensional space, specifically configured with n_components=40n\_components=40, n_neighbors=120n\_neighbors=120, and min_dist=0.08min\_dist=0.08 (with a fixed random_state=42random\_state=42 for reproducibility). UMAP utilized the cosine metric, which aligns with the semantic nature of the embeddings. Clustering was performed using HDBSCAN, which is optimized for identifying clusters of varying densities; the algorithm was set with a significantly smaller min_cluster_size=50min\_cluster\_size=50 and min_samples=1min\_samples=1, employing the eom (Excess of Mass) cluster selection method. Topic representation was established using C-TF-IDF, configured to filter vocabulary using $\text{MAX\_DF\_PERCENT}=0.9$ and a $\text{MIN\_DF\_ABS\_FLOOR}=50$, utilizing an n-gram range of (1, 2). The full hierarchical structure, derived from the HDBSCAN tree and refined by an external LLM (ChatGPT-5) for descriptive naming, resulted in the fine-grained multi-level taxonomy of over 300 topics presented in Appendix˜C.
Table A2: Sample list of fine-grained topics extracted from ResearchDBResearchDB, illustrating the specificity and coverage of the taxonomy.
@@
Core ML 042 Causal Inference for Robustness and Generalization do-calculus, invariant prediction, domain generalization
@@
Hierarchical Topic Tree Structure. A key advantage of the HDBSCAN algorithm is its inherent ability to produce a Hierarchical Structure. We leverage this structure as an organizational framework, providing a non-strict view of clustering relationships above the 324 fine-grained topics. This tree structure is the basis for our hierarchical retrieval system, allowing users to navigate from macro-level concepts (near the root) down to specific fine-grained topics (at the leaves), as shown in Fig.˜A1.
@@
Topic Naming and Examples. The naming of each fine-grained topic underwent strict quality control: C-TF-IDF was used to extract representative keywords, which were then analyzed and refined by ChatGPT-5 based on paper abstracts within the cluster. This process generated highly semantically clear Topic Names (e.g., 140_Approximation and Fair Clustering Algorithms). Every paper in $ResearchDB$ is uniquely tagged with such a topic name. Tab.˜A2 below presents a sample of the most representative fine-grained topics. We also have selected topics with ids ranging from 1 to 20 and listed them as Tab.˜A3
Table A3: Twenty Sample Fine-Grained Topics from ResearchDBResearchDB, Illustrating the Taxonomy’s Scope and Detail.
@@
Evaluation Results. Overall validation demonstrates that the $ResearchDB$ retrieval system, driven by multidimensional knowledge profiling, significantly outperforms traditional keyword or embedding-only retrieval baselines on complex, multi-faceted queries. To illustrate this advantage, we conducted a qualitative assessment contrasting the $ResearchDB$ output with a ChatGPT-5 baseline on a complex query: “How to achieve real-time, high-fidelity speech-to-gesture generation using decoupled diffusion models.” The $ResearchDB$-supported answer (akin to an Academic Review) showed clear superiority in Fig.˜A2:
@@
Efficiency. To evaluate the scalability of ResearchDB, we benchmarked the knowledge profiling pipeline on a cluster equipped with 8 NVIDIA A100 (80GB) GPUs.
  • Knowledge Extraction Throughput: By leveraging a distributed parallel processing framework with vLLM for inference acceleration, the system achieved a steady-state throughput of approximately 870 full papers per hour. The entire corpus of 104,482 papers was processed in 120 hours (5 days) of wall-clock time, totaling 960 GPU hours.
  • Data Sharing and Accessibility: To facilitate future research and ensure reproducibility, we will release the fully structured ResearchDB, containing multidimensional profiles (e.g., methods, datasets, and limitations) for all 104,482 papers. This pre-processed knowledge base allows researchers to perform large-scale trend analysis and complex semantic queries without incurring the high computational costs of initial parsing.
  • Query Response: For the end-user, the parallelized intent recognition module decomposes complex queries into sub-tasks that are executed concurrently, maintaining an average end-to-end response time of $\sim$1.2 seconds.